Hello please assist me with this question.

1. Let the function f be defined by y = f(x), where x and f(x) are real numbers. Find f(2), f(−7), f(k), and f(k2 − 1). f(x) = 3 x − 5

2. Evaluate the expression. P(12, 2) · C(12, 2)

3. Find an equation in standard form of the parabola described.

Vertex (8, 3); focus at (8, 10)

4. The population of Eagle River is growing exponentially according to the model P(t) = 475(1.3)t

people, where t is measured in years from the present date. Find the population in 3 years. (Round your answer to the nearest whole number.)

I had the answer 1043, which is incorrect.
Answer #1

f(x) 32-5 3(2)-5 3(-7) -S b) f(-7)= ay f(2)= -21-5 6-5 26 d) f(x-1)= 3(K-1) -5 3k-s ♡ f(K) ✓ = 3k?. 3-5 3k² 8 ✓